OpenStudy (anonymous):

A collection of 70 coins consisting of dimes, quarters, and half-dollars has a value of 17.75. there are three timesas many quarters as dimes. Find the number of ecah kind of coin

OpenStudy (saifoo.khan):

x+y+z=70;x=3y;4y+z=70 .25(3y)+.10y+.5z=17.75 .85y+.5z=17.75 85y+50z=1775 200y+50z=3500 ---------------- 115y=1725 y=15 x=45 z=10 There are 45 quarters, 15 dimes, and 10 half-dollars in the coin collection..

OpenStudy (anonymous):

Thank you, but i still don't understand how you got the result. Would you please explain with a little bit of details. Thanks oce again :)